# Is Your AI Truly Yours? Leveraging Blockchain for Copyrights, Provenance, and Lineage

### v1

@@ IV-B Functional Operations
First, we need to locate the corresponding LVC smart contract for validating the license. This can be accomplished using another hash map where the license type $\mathsf{typeID}$ serves as the key and LVC’s address as the value. Consequently, this lookup operation can be performed in constant time, i.e., $O(1)$. Next, we need to invoke the identified LVC contract to determine the license validity. The time complexity of the LVC contract is directly proportional to the number of environment variables that need validation. We abstract this time complexity as $O(|E|)$, where E is the set of environment variables to validate. Therefore, the overall time complexity is $O(1)+O(|E|)=O(|E|)$.
Obtain licensed datasets. The $\mathsf{getLicensedDatasets}$ operation retrieves the list of dataset identifiers $\mathsf{datasetId}$s each with a valid license. It entails executing $getDatasetLicense$ and $checkLicenseValidity$ operations for each dataset. Given D datasets, the overall time complexity is $O(|D|\times\{O(1)+O(E)\}=O(|D||E|)$.
Obtain authorized datasets by license. Given a license identifier $\mathsf{licenseId}$, the $\mathsf{getDatasetsByLicense}$ operation retrieves datasets covered by the license. This operation performs a search of the DMR using the licenseId, resulting in a time complexity of $O(1)$.
Obtain authorized models by license. Given a license identifier $\mathsf{licenseId}$, the $\mathsf{getModelsByLicense}$ operation retrieves the metadata of the models covered by the license. This operation entails executing $\mathsf{getDatasetsByLicense}$, followed by conducting a graph traversal that goes from each dataset to the models trained on it (including child models indirectly trained on it). In the worst case, the overall time complexity is $O(|D||M|)$.
Obtain model datasets. Given a model identifier $\mathsf{modelId}$, the $\mathsf{getModelDatasets}$ operation retrieves the identifiers of its training datasets. This operation extracts the $\mathsf{datasetList}$ attribute from the provided model and its upstream source models. If the provided model has a set of M upstream models, then the time complexity is $O(|M|)$.
IV-C License Renewal
@@
IV-C1 License Renewal Check
The framework supports three types of license renewal checks (LRCs): license-driven, dataset-driven, and model-driven.
@@ IV-C2 License Renewal
A license renewal involves adding a new bilaterally signed license to the license registry, rather than updating existing records. This enables AOs and COs to access all historical licenses to prove regulatory compliance and avoid any disputes. After a new license has been added, an AO can execute the $\mathsf{getModelsByLicense}$ operation to gather the identifiers of datasets and models that depend on the renewed license. Then the DMR records of datasets are updated to reference the new license. As the list of dependent datasets and models can now be considered eligible for training, their identifiers are also removed from the blacklist.
IV-D Operation Atomicity
@@
It is observed that several stages (i.e., S.1, S.3, and License Renewal) involve the update of multiple records. Apart from ensuring integrity and immutability, another advantage of maintaining DMR, license registry, and MMR on-chain is that such multiple updates are guaranteed to be atomic. This is because smart contracts can ensure atomicity where the actions included in one transaction either all take effect or none of them take effect. Therefore, care needs to be taken during implementation to ensure that all updates within a stage should be included in the same transaction.
V Implementation on Daml
V-A Blockchain Platform
We implemented the proof of concept framework using Daml (Digital Asset Modeling Language) [27] atop the Canton blockchain ledger protocol [28]. We used Daml version 2.8.3 with the corresponding Daml ledger model and Canton protocol.
The rationale for utilizing Daml in implementing our framework is twofold.
First, in Daml, a smart contract codifies the terms of the agreement between parties, including the rights and obligations of each party. A Daml contract template describes the data schema of the contract and rules for manipulating the data. In our implementation, these contract templates align closely with the data models outlined in Sec.IV-A.
Second, the Daml ledger model uses smart contracts as privacy-enabled data containers, enforcing data access controls as specified in the contract. The access control permissions of a party/user are explicitly stated in contract templates by assigning the party one of the following predefined roles:
  • signatory: A party whose authority is required to create the contract or archive it. Every contract must have at least one signatory. Signatories are guaranteed to see actions on that contract.
@@ V-A Blockchain Platform
While similar approaches can be implemented with other smart contract languages, the contract-level access control typically does not extend to the ledger itself. For example, although chaincodes in Hyperledger Fabric can enforce access control, all channel members still synchronize the entire ledger [50]. In contrast, the Canton blockchain ledger protocol extends the Daml ledger model to the ledger level, where a blockchain node synchronizes only the contract data relevant to its party permissions. Therefore, by adopting Daml and Canton, we facilitate the integration of many AOs and COs onto the same IBis platform, while ensuring that commercially sensitive license agreements between an AO and COs, dataset metadata, and model metadata remain concealed from other parties, even at the ledger level. Table III lists the resources in IBis along with their authorizations.
@@
DMR AOs Datasets added by the actor.
License Registry AOs/COs Licenses signed by the actor.
MMR AOs Models owned by the actor.
LVC API AOs/COs Licenses signed by the actor.
CLM API COs Licenses drafted by the actor.
V-B Smart Contract Implementation
Listing 1 illustrates the license smart contract template, constructed based on the data schema in Table I. The $\mathsf{copyrightOwner}$ and $\mathsf{modelOwner}$ are designated as signatories of the template (Line 10). This reflects the bilateral nature of the license agreement. Consequently, authorization from both parties is required to create a license contract. Subsequently, once the contract is created, both parties can access it, while no other party has access to this contract either on-chain or on-ledger. Hence, this implementation closely aligns with the access control requirements specified in Table III. In addition, the identifier attribute of each license serves as the primary key (specified by Line 12-13), which can facilitate efficient queries during the graph traversal.
@@ V-B Smart Contract Implementation
Listing 1: Daml smart contract template for license.
@@
V-C Multi-party Signing
@@
Listing 2: Daml smart contract template for license agreement.
In our implementation, the proposal contract takes the form of the draft license agreement, and it is depicted in Listing 2. $\mathsf{LicenseAgreement}$ shares the same data schema as $\mathsf{License}$ (see Listing 1), with the $\mathsf{copyrightOwner}$ designated as the signatory. This ensure only the $\mathsf{copyrightOwner}$ has the authority to create a contract. Alternatively, $\mathsf{modelOwner}$, as the controller of the $\mathsf{Accept}$ choice in Line 16, has the authority to exercise the choice, resulting in the creation of a $\mathsf{License}$ contract. This setup ensures that only the $\mathsf{modelOwner}$ specified in a $\mathsf{LicenseAgreement}$ contract has the authority to exercise the $\mathsf{Accept}$ choice of that contract.
@@ VI Evaluation
VI-A Experimental Setup
Our proof-of-concept IBis implementation was deployed on a private Canton blockchain comprising three nodes. All nodes were hosted on the same AWS EC2 $\mathsf{t2.xlarge}$ instance with four virtual CPUs and 16GB of RAM. While Daml provides a range of options for data storage, PostgresSQL, running within Docker containers, is chosen as the data storage to persist node data. Our source code of the performance test is available [^6] Testing script: https://github.com/yilin-sai/ai-copyright-framework.
@@ VI-A Experimental Setup
VI-B Experimental Parameters
In real-world scenarios, the framework hosts data, including dataset metadata, license, and model metadata, contributed by various AOs and COs. These stakeholders engage in executing functional operations (outlined in Table II) to realize data provenance and copyright management. Ensuring the efficiency of operations, particularly those involving complex graph traversals, is paramount in this context. The experimental environment is set up the following parameters:
  • The framework accommodates $N$ AOs, where each scrape $D$ datasets for model training. Therefore, the framework host a total of $N\times D$ datasets.
@@ VI-B Experimental Parameters
Note that there are five parameters during the experimental setup, namely $N$, $D$, $L$, $M$, and $T$. The system workload can be scaled up by increasing these parameters. In our evaluation, adhering to the control variates method, we measure the performance by varying each parameter individually while keeping the other parameters fixed. Table IV lists the values of the four parameters that remain fixed while adjusting the remaining parameter.
@@
N 10 to 100 10 10 10 10
D 10 10 to 100 10 10 100
L 10 10 10 to 100 10 10
M 1 1 1 1 to 10 1
T 10 10 10 10 10 to 100
VI-C Evaluation of Fetching Model Licenses
Fig.8 illustrates the variations in execution time corresponding to incremental adjustments in each of the five parameters. As explained above, each data point represents the average execution time of $\mathsf{getModelLicenses}$ in ten executions, with error bars indicating the standard deviation. The result reveals that the execution time of $\mathsf{getModelLicenses}$ increases linearly with the augmentation of the number of models in the model chain $M$ and number of training datasets of each model $T$. This correlation is logical, as elevating $M$ augments the model training graph depth (see Fig.6), whereas elevating $T$ expands its breadth.
@@ VI-C Evaluation of Fetching Model Licenses
The outcomes also indicate that the values of the number of scraped datasets per model owner $D$, model owners $N$, and licenses per model owner $L$ exert no discernible influence on the performance of $\mathsf{getModelLicenses}$. In theory, these three parameters do not impact the graph size; hence they have negligible effect on performance. However, theoretically, they could affect performance as querying a record using its identifier might slow down with a greater number of records. However, our optimization efforts, such as designating data, model, and license identifiers as the primary key of a record (cf. Sec.V-B), mitigate any observable impact of increased record numbers. The results demonstrate that the performance of $\mathsf{getModelLicenses}$ operation remains consistent regardless of the number of model owners in the system, datasets they scrape, or licenses they acquire, thereby affirming the scalability of the operation.
@@
Figure 8: Performance of fetching model licenses.
VI-D Evaluation of Fetching Model Datasets
Fig.9 illustrates the variations in execution time corresponding to incremental adjustments in each of the five parameters. The operation can be viewed as a sub-operation of $\mathsf{getModelLicenses}$ that traverses the entire graph from a model to licenses. In contrast, the $\mathsf{getModelDatasets}$ operation stops the traversal early at the level of datasets. Therefore, the two operations share many common characteristics in terms of performance. The results highlight a notable correlation between the execution time and the number of models in the model chain $M$ and training datasets of each model $T$. As $M$ increases, indicating a deeper model-data graph structure, and $T$ expands, indicating a broader breadth of the graph, the execution time rises linearly. This relationship stems from the increased computational complexity associated with traversing deeper and wider graphs.
Moreover, experiments show that variations in the number of scraped datasets per model owner $D$, model owners $N$, and licenses per model owner $L$ do not significantly impact performance. This observation aligns with similar findings for $\mathsf{getModelLicenses}$ and underscores the operation’s scalability. The evaluated operation consistently maintains its performance regardless of the number of model owners in the system, datasets they scrape, or licenses they acquire, reflecting the scalability and efficiency in managing data provenance.
@@ VI-D Evaluation of Fetching Model Datasets
Figure 9: Performance of fetching model datasets.
@@
Figure 10: Performance of fetching authorized models.
VI-E Evaluation of Fetching Authorized Models
Fig.10 illustrates the variations in execution time corresponding to incremental adjustments in each of the five parameters. The result exhibits an overall increasing trend in execution time with the increasing number of models in the model chain $M$. This observation is intuitive, as the operation necessitates traversing more models as the chain of related model lengthens. However, the average performance displays oscillations as $M$ increases, accompanied by high standard deviations for each data point. These fluctuations and high standard deviations stem from the presence of redundancy of the licenses and datasets.
In real-world scenarios, redundancy often occurs because licenses may be acquired in advance, before the corresponding data is scraped, or datasets may be stored without immediate model training. Our experimental setup reflects these real-world complexities, resulting in some executions being faster due to the operation encountering redundant licenses or datasets. Consequently, the graph traversal terminates early in these instances, leading to variations in execution times. This phenomenon also explains the observed high standard deviations in the other charts.
The results also indicate that the execution time remains constant as the number of training datasets per model $T$ increases. This is because increasing $T$ does not impact the size of the graph starting from a particular license. However, the execution time linearly increases with the increasing number of scraped datasets per model owner $D$. This phenomenon occurs because as more datasets become associated with a license, the graph starting from that license experiences an increase in breadth, consequently prolonging the traversal time.
Additionally, the performance remains consistent with the increasing number of model owners $N$ and licenses per model owner $L$, mirroring the behavior observed in the previous operations. This consistency affirms the scalability of IBis to accommodate a large number of users and licenses.
VI-F Discussions Between Evaluated Operations

### v2

@@ IV-C Validating Licenses
First, we need to locate the corresponding LVC smart contract for validating the license. This can be accomplished using another hash map where the license type $\mathsf{typeID}$ serves as the key and LVC’s address as the value. Consequently, this lookup operation can be performed in constant time, i.e., $O(1)$. Next, we need to invoke the identified LVC contract to determine the license validity. The time complexity of the LVC contract is directly proportional to the number of environment variables that need validation. We abstract this time complexity as $O(|K|)$, where K is the set of environment variables to validate. The overall time complexity is $O(1)+O(|K|)=O(|K|)$.
Here, we give an example (Listing 1). This LVC template is triggered by the $\mathsf{checkLicenseValidity}$ operation and verifies that the license is currently active based on temporal conditions. The modular design of LVCs allows for extensible contract templates supporting various types of constraints (e.g., geographic, exclusivity, usage count), which can be selected based on the license’s $\mathsf{typeId}$.
1template TimeBoundLVC with
2 licenseId: Text
3 validFrom: Time
4 validUntil: Optional Time
5 currentTime: Time
6 where
7 signatory licenseVerifier
8
9 choice CheckValidity : Bool
10 controller licenseVerifier
11 do
12 let isValidStart = currentTime >= validFrom
13 let isValidEnd = case validUntil of
14 Some end -> currentTime <= end
15 None -> True
16 return (isValidStart && isValidEnd)
Listing 1: Example LVC smart contract template for validating a time-bound license.
IV-D Obtaining Data and Models via Licenses
Obtain licensed datasets. The $\mathsf{getLicensedDatasets}$ operation retrieves the list of dataset identifiers $\mathsf{datasetId}$s each with a valid license. It entails executing $\mathsf{getDatasetLicense}$ and $\mathsf{checkLicenseValidity}$ operations for each dataset. Given D datasets, the overall time complexity is $O(|D|\times\{O(1)+O(K)\}=O(|D||K|)$.
Obtain authorized datasets by license. Given a license identifier $\mathsf{licenseId}$, the $\mathsf{getDatasetsByLicense}$ operation retrieves datasets covered by the license. This operation performs a search of the DMR using the $\mathsf{licenseId}$, resulting in a time complexity of $O(1)$.
Obtain authorized models by license. Given a license identifier $\mathsf{licenseId}$, the $\mathsf{getModelsByLicense}$ operation retrieves the metadata of the models covered by the license. This operation begins by executing $\mathsf{getDatasetsByLicense}$, which identifies all datasets associated with the given license. For each of these datasets, the operation performs a graph traversal to retrieve the models directly trained on them and any child models indirectly connected through shared datasets. The overall time complexity consists of two components: $O(|D|)$ for identifying the datasets linked to the license and $O(|M||T|)$ for traversing the graph to retrieve the connected models and their relationships. Consequently, the total complexity is $O(D+|M||T|)$, where $D$ is the total number of datasets, $M$ is the total number of models, and $T$ is the average number of datasets per model.
Obtain model datasets. Given a model identifier $\mathsf{modelId}$, the $\mathsf{getModelDatasets}$ operation retrieves the identifiers of its training datasets. During the first traversal or initialization, this operation iterates through the provided model and its upstream source models, retrieving the associated datasets. If the model has a set of M upstream models, each linked to T datasets, the time complexity is $O(|M||T|)$ for the initialization phase.
IV-E Renewing License
@@
License renewal check. The framework supports three types of license renewal checks (LRCs): license-driven, dataset-driven, and model-driven.
@@
License renewal. A license renewal involves adding a new bilaterally signed license to the license registry, rather than updating existing records. This enables AOs and COs to access all historical licenses to prove regulatory compliance and avoid any disputes. After a new license has been added, an AO can execute the $\mathsf{getModelsByLicense}$ operation to gather the identifiers of datasets and models that depend on the renewed license. Then the DMR records of datasets are updated to reference the new license. As the list of dependent datasets and models can now be considered eligible for training, their identifiers are also removed from the blacklist.
IV-F More Operational Features
@@
Operation atomicity. It is observed that several stages (i.e., S.1, S.3, and License Renewal) involve the update of multiple records. Apart from ensuring integrity and immutability, another advantage of maintaining DMR, license registry, and MMR on-chain is that such multiple updates are guaranteed to be atomic. This is because smart contracts can ensure atomicity where the actions included in one transaction either all take effect or none of them take effect. Therefore, care needs to be taken during implementation to ensure that all updates within a stage should be included in the same transaction.
License linkage. In IBis, each dataset is assigned a license at the time of registration through the $\mathsf{licenseId}$ attribute. This linkage is immutable and ensures that all subsequent uses of the dataset are bound by the same licensing terms. The license itself enumerates the covered datasets via its $\mathsf{datasetList}$ attribute. This bidirectional mapping supports consistent enforcement of usage policies and enables automatic validation of licensing conditions during model training and retraining. The use of type-specific LVC contracts (see Sec. IV-B) ensures that the correct license validation logic is applied in each case.
Dataset reuse. When a dataset is reused across multiple models, original licensing terms remain applicable. The reuse is authorized as long as the license associated with the dataset remains valid. During model creation or retraining, the system verifies license validity through $\mathsf{checkLicenseValidity}$ and ensures that all datasets listed in the model’s $\mathsf{datasetList}$ are covered by licenses returned by $\mathsf{getModelLicenses}$. Since licenses may impose usage constraints (e.g., valid period, geographic restrictions), this step guarantees ongoing legal compliance for every instance of dataset reuse.
Traceability. To support auditing and provenance, IBis explicitly records the lineage of models and datasets. Each dataset maintains a $\mathsf{modelList}$ tracking all models trained on it, while each model retains a $\mathsf{datasetList}$ identifying its training inputs. For retrained models, the $\mathsf{sourceModelId}$ and $\mathsf{childModelList}$ attributes form a directed acyclic graph (DAG), capturing the complete history of derivations. These structural links are embedded in the smart contract state and are queryable via the $\mathsf{getModelDatasets}$ and $\mathsf{getModelsByLicense}$ operations. As a result, IBis provides fine-grained, tamper-proof traceability across both horizontal (multi-model) and vertical (model-chain) reuse scenarios.
V Implementation
V-A Overview
Prototype stack. We deploy Canton [39] network (v2.8.3) with three controlled nodes on AWS $\mathsf{t2.xlarge}$ instances (configuration by 4vCPU, 16GB RAM, Ubuntu22.04). Each node runs a Participant+Synchroniser pair backed by Docker-ised PostgreSQL15.
Component map. Fig.7 links the on-chain contracts (DMR, MMR, License) to off-chain services (MSS, CLM); Table III lists party-centric access scopes enforced by Canton.
Code footprint. Core Daml templates $\approx 380$ LOC, HTTP/gRPC glue $\approx 210$ LOC, evaluation harness $\approx 150$ LOC.
Reproducibility. We wrap up experiments and a single command line (i.e., \$ $\mathsf{docker-composeup}$) can instantiate the full network and re-run Experiments 1–4; scripts are open-sourced at https://github.com/yilin-sai/ai-copyright-framework.
V-B Blockchain Platform
We implement IBis in Daml (Digital Asset Modelling Language) atop the Canton ledger protocol.
Contract-as-data model. Daml templates codify both schema and behaviour. We can map the Dataset/Model/License records of Table II to verifiable contracts with one-to-one clarity.
Ledger-level privacy. Canton extends Daml’s role system (signatory, controller, observer, choice-observer) so each node synchronises only those contracts it is authorised to see; commercial terms inside a License stay invisible to third parties even on-ledger.
@@ V-B Blockchain Platform
Why not Fabric / generic EVM? Hyperledger Fabric chaincode can enforce ACLs but still replicates the entire channel ledger to all members [60]. Public-chain EVM contracts disclose everything. Canton’s party-centric sync therefore offers the best trade-off between immutability and confidentiality for multi-tenant AI marketplaces.
Therefore, by adopting Daml and Canton, we facilitate the integration of many AOs and COs onto the same IBis platform, while ensuring that commercially sensitive license agreements between an AO and COs, dataset metadata, and model metadata remain concealed from other parties, even at the ledger level. Table III lists the resources in IBis along with their authorizations.
@@
DMR AOs Datasets added by the actor
License Registry AOs/COs Licenses signed by the actor
MMR AOs Models owned by the actor
LVC API AOs/COs Licenses signed by the actor
CLM API COs Licenses drafted by the actor
V-C Smart-Contract Implementation
Listing 2 illustrates the license smart contract template, constructed based on the data schema in Table II. The $\mathsf{copyrightOwner}$ and $\mathsf{modelOwner}$ are designated as signatories of the template (Line 10). This reflects the bilateral nature of the license agreement. Consequently, authorization from both parties is required to create a license contract. Subsequently, once the contract is created, both parties can access it, while no other party has access to this contract either on-chain or on-ledger. Hence, this implementation closely aligns with the access control requirements specified in Table III. In addition, the identifier attribute of each license serves as the primary key (specified by Line 12-13), which can facilitate efficient queries during the graph traversal.
@@ V-C Smart-Contract Implementation
Listing 2: Daml smart contract template for license.
@@
V-D Multi-party Signing
@@
In our implementation, the proposal contract takes the form of the draft license agreement, and it is depicted in Listing 3. $\mathsf{LicenseAgreement}$ shares the same data schema as $\mathsf{License}$ (see Listing 2), with the $\mathsf{copyrightOwner}$ designated as the signatory. This ensure only the $\mathsf{copyrightOwner}$ has the authority to create a contract. Alternatively, $\mathsf{modelOwner}$, as the controller of the $\mathsf{Accept}$ choice in Line 16, has the authority to exercise the choice, resulting in the creation of a $\mathsf{License}$ contract. This setup ensures that only the $\mathsf{modelOwner}$ specified in a $\mathsf{LicenseAgreement}$ contract has the authority to exercise the $\mathsf{Accept}$ choice of that contract.
@@
Listing 3: Daml smart contract template for license agreement.
V-E Operational Footprint
Gas / storage. A complete write–read cycle touches four contracts and stores $\approx 3.5$kB; even with $10^{3}$ datasets/models/licenses the ledger is $\leq$ 15MB.
Performance hooks. Primary-key indices and batched Canton commits keep the $P95$ end-to-end latency below 1.2 s at 1 000 req/s (§VI) without heavyweight zk-proofs or HE / sMPC.
@@ VI Evaluation
Our proof-of-concept IBis implementation was deployed on a private Canton blockchain comprising three nodes. All nodes were hosted on the same AWS EC2 $\mathsf{t2.xlarge}$ instance with four virtual CPUs and 16GB of RAM. While Daml provides a range of options for data storage, PostgresSQL, running within Docker containers, is chosen as the data storage to persist node data. Our source code of the performance test is available [^6] Testing script: https://github.com/yilin-sai/ai-copyright-framework. In the following sections, we present four sets of experiments: (1) baseline evaluation of core search operations, (2) comparative analysis of execution overhead against other blockchain-based designs, (3) scalability under varying load and storage conditions, and (4) robustness against adversarial attacks targeting provenance and signature integrity.
VI-A Experiment-1: Searching Operations
@@
Benchmarks and baselines. In the absence of a unified benchmark for evaluating methods in this context, we focus on execution time as the primary evaluation criterion. This metric is reasonable and sufficient given that our method, IBis, is implemented on a blockchain and emphasizes efficient searching when providing copyright protection for models and datasets. Moreover, the focus on execution time aligns with common practices in evaluating non-functional requirements, especially for blockchain-based software platforms [61, 62, 63], as functional requirements can vary significantly and are often implemented via heterogeneous protocol designs, making precise cross-method evaluations challenging.
To highlight the advantages of IBis, we compare it against two widely used architectural patterns: Flat/Sequential Search and Full Graph Search. These baselines represent common designs in real-world systems such as e-Commerce [62] and e-Healthcare [64], and also subsume recent blockchain-based copyright systems [65, 66, 67].
  • Flat/Sequential Search [62, 65, 66]: Assets are treated as isolated records with no structured relationships. Systems of this type store either individual entries (e.g., CIDs and ownership metadata [65]) or simple one-way chains per asset [66]. Each query scans a linear list or full dataset to validate, retrieve, or compare records. For example, $\mathsf{getDatasetLicense}$ requires $O(|D|)$ time compared to $O(1)$ in IBis; $\mathsf{getModelLicenses}$ is $O(|D||M|)$ vs. $O(|M||T|)$, and $\mathsf{getModelsByLicense}$ is $O(|D||M|)$ vs. $O(|D|+|M||T|)$. These access costs grow rapidly with scale and offer no built-in mechanism for capturing cross-asset dependencies.
  • Full Graph Search [64, 67]: More expressive systems represent licenses, evidence, and ownership as interconnected graph structures. For instance, [67] uses RDF triples and semantic links to connect works, claims, and complaints, with queries executed via external graph indexers. However, without pruning or indexing, operations such as $\mathsf{getModelLicenses}$ must walk all linked paths, resulting in $O(|M||T|+|D|)$ complexity compared to $O(|M||T|)$ in IBis. These systems capture richer semantics but suffer from high traversal costs and limited performance tuning.
IBis generalizes and improves upon both categories. It supports rich inter-asset semantics like the graph-based approaches while maintaining low-complexity queries through DAG indexing and caching. Because our design operates at the indexing layer, it can be added atop existing ledger infrastructures without changing their consensus or storage logic. This allows flat or graph-based systems to adopt IBis for efficient, provenance-aware queries and recursive license value flows.
Experimental parameters. In real-world scenarios, the framework hosts data, including dataset metadata, license, and model metadata, contributed by various AOs and COs. These stakeholders engage in executing functional operations (outlined in Table V) to realize data provenance and copyright management. Ensuring the efficiency of operations, particularly those involving complex graph traversals, is paramount in this context. The experimental environment is set up the following parameters:
  • The framework accommodates $N$ AOs, where each scrapes $D$ datasets for model training. Therefore, the framework host a total of $N\times D$ datasets.
@@
There are five parameters during the experimental setup, namely $N$, $D$, $L$, $M$, and $T$. The system workload can be scaled up by increasing these parameters. In our evaluation, adhering to the control variates method, we measure the performance by varying each parameter individually while keeping the other parameters fixed. Table IV lists the values of the four parameters that remain fixed while adjusting the remaining parameter.
@@
N 10 to 100 10 10 1 10
D 10 10 to 100 10 1 10
L 10 10 10 to 100 1 10
M 10 10 10 1 to 10 10
T 10 100 10 1 10 to 100
Extreme cases. While Table IV lists typical parameter ranges for evaluation, we intentionally do not explore extreme cases such as $M\gg T$, $M\ll T$, or very large $D$. These conditions are uncommon in practical AI workflows. In most scenarios, the average model retraining depth $M$ ranges between $1$ and $10$, while the number of training datasets $T$ remains within $10$ to $100$. To prevent unbounded graph growth and ensure runtime stability under rare edge cases, IBis incorporates bounding strategies such as retraining checkpoints, anchor nodes, and lineage aggregation. These allow IBis to maintain traversal efficiency and theoretical consistency even under potentially adversarial graph topologies. In all tested regimes, performance trends align with the expected complexities outlined in Table V.
Complexity analysis. We provide the complexity analysis. Table V lists the time complexity of operations and the entities authorized to perform them. We assume that the on-chain license registry, DMR, and MMR are implemented as hash maps on a smart contract, resulting in a time complexity of $O(1)$ for searching them.
TABLE V: Operations, complexities, and authorizations.
Operation Complexity Authorisations
$\mathsf{getDatasetLicense}$ $O(1)$ AOs
$\mathsf{getModelLicenses}$ $O(|M||T|)$ AOs
$\mathsf{checkLicenseValidity}$ $O(|K|)$ AOs/COs
$\mathsf{getLicensedDatasets}$ $O(|D||K|)$ AOs
$\mathsf{getDatasetsByLicense}$ $O(1)$ AOs
$\mathsf{getModelsByLicense}$ $O(|D|+|M||T|)$ AOs
$\mathsf{getModelDatasets}$ $O(|M||T|)$ AOs
Evaluation of fetching model licenses. Fig.8 illustrates the variations in execution time for the $\mathsf{getModelLicenses}$ operation. Each data point represents the average execution time over 10 executions, with error bars indicating the standard deviation.
The execution time increases linearly with the number of models in the model chain ($M$) for all three methods, consistent with their theoretical complexities (cf. Fig.8). IBis demonstrates the best performance due to its targeted graph traversal, which efficiently limits operations to $T$, the datasets directly linked to the models, resulting in a complexity of $O(MT)$. In contrast, Flat Search performs the worst, with a complexity of $O(MD)$ that involves scanning all $D$ datasets for every model, causing significant overhead for large $D$. Full Graph Search performs better than Flat Search, benefiting from $O(MT)$ traversal for relevant datasets, but its additional $O(D)$ term for scanning all datasets for dataset-to-license mappings creates noticeable overhead, especially as $D$ increases.
A similar trend is observed with the number of training datasets per model ($T$), where execution time grows linearly for both IBis and Full Graph Search (cf. Fig.8). The targeted traversal in IBis ensures that only relevant datasets are processed, maintaining its efficiency as $T$ increases. However, Full Graph Search incurs additional costs due to its lack of explicit dataset-to-license mappings. Flat Search remains unaffected by $T$, as its execution time is dominated by $D=100$, rather than the specific number of datasets linked to each model. The effect of increasing the total number of scraped datasets per model owner ($D$) is most pronounced in Flat Search, where execution time grows steeply due to its dependency on $D$ in the $O(MD)$ complexity (cf. Fig.8). Full Graph Search also exhibits growth with increasing $D$, as it scans all datasets during the dataset-to-license mapping phase. In contrast, IBis is largely unaffected by the increase in $D$, as it relies on targeted traversal of $T$ datasets, bypassing irrelevant entries. This scalability highlights the advantage of IBis for systems with large datasets. The number of model owners ($N$) and licenses per model owner ($L$) have negligible impact on execution time for all three methods (cf. Figs.88). This is consistent with theoretical expectations, as $N$ and $L$ do not influence the graph size or traversal complexity. Furthermore, the use of optimizations, such as primary keys for identifiers (Sec.V-C), ensures that querying operations remain efficient regardless of the number of records in the system.
@@
The outcomes also indicate that the values of the number of scraped datasets per model owner $D$, model owners $N$, and licenses per model owner $L$ exert no discernible influence on the performance of $\mathsf{getModelLicenses}$. In theory, these three parameters do not impact the graph size; hence they have negligible effect on performance. However, theoretically, they could affect performance as querying a record using its identifier might slow down with a greater number of records. However, our optimization efforts, such as designating data, model, and license identifiers as the primary key of a record (cf. Sec.V-C), mitigate any observable impact of increased record numbers. The results demonstrate that the performance of $\mathsf{getModelLicenses}$ operation remains consistent regardless of the number of model owners in the system, datasets they scrape, or licenses they acquire, thereby affirming the scalability of the operation.
Evaluation of fetching model datasets. Fig.9 presents the execution time of the $\mathsf{getModelDatasets}$ operation under varying parameters. The results align with the theoretical complexities and highlight clear differences in performance as the parameters scale. As the number of models in the model chain ($M$) increases, the execution time grows linearly for all methods. Both IBis and Full Graph Search demonstrate similar performance, consistent with their shared complexity of $O(MT)$, as they efficiently traverse only the $T$ datasets associated with each model (cf. Fig.9). Flat Search, however, performs significantly worse, reflecting its $O(MD)$ complexity, which involves processing all $D$ datasets for each model. This overhead becomes more pronounced as $D$ increases.
Execution time also scales linearly with the number of training datasets per model ($T$) for IBis and Full Graph Search, as their traversal depends on the number of datasets directly linked to the models (cf. Fig.9). Flat Search, by contrast, remains unaffected by changes in $T$, since its execution time is determined by $D$ and $M$. This behavior underscores the inefficiency of Flat Search in targeting relevant datasets compared to the more focused approaches of IBis and Full Graph Search.
When varying the number of scraped datasets per model owner ($D$), the number of model owners ($N$), and the number of licenses per model owner ($L$), the results reveal distinct trends (cf. Figs.99). Larger $D$ impacts the performance of Full Graph Search, as its $O(MT+D)$ complexity includes scanning all datasets, leading to a steady increase in execution time. In contrast, IBis maintains stable performance by limiting operations to $T$ datasets, bypassing the irrelevant portions of $D$. Flat Search shows relatively constant execution times with increasing $D$, as it processes all $D$ datasets regardless of their relevance, making the increase less noticeable. For $N$ and $L$, the execution times remain relatively stable across all methods, as these parameters do not directly influence graph traversal. The consistent performance of IBis in all cases validates its efficiency and scalability compared to Flat Search and Full Graph Search.
@@
Figure 8: Comparison between IBis and baselines in terms of performance of fetching model licenses.
@@
Figure 9: Comparison between IBis and baselines in terms of performance of fetching model datasets.
@@
Figure 10: Comparison between IBis and baselines in terms of performance of fetching authorized models.
Evaluation of fetching authorized models. Fig.10 illustrates the performance of the $\mathsf{getModelsByLicense}$ operation under varying system parameters. The execution time trends observed validate the theoretical complexities associated with each method. In Fig.10, as the number of models in the model chain ($M$) increases, both IBis and Full Graph Search exhibit a similar linear growth in execution time, consistent with their shared complexity of $O(MT+D)$. The efficient traversal of only the $T$ datasets linked to the models, combined with a limited dependency on $D$, explains their comparable performance. Flat Search, on the other hand, incurs similar execution times in this experiment due to $T\simeq D$ in the experimental setup, aligning with its $O(MD)$ complexity. However, in scenarios where $T\ll D$, Flat Search would exhibit significantly higher execution times due to its inefficiency in processing all $D$ datasets for each model.
As the number of training datasets per model ($T$) increases, the execution time grows linearly for both IBis and Full Graph Search (cf. Fig.10). This behavior aligns with their $O(MT)$ complexity, as the traversal scales with the number of datasets connected to the models. In contrast, Flat Search remains unaffected by changes in $T$, as its performance depends solely on $M$ and $D$ (with $D$ fixed at 100 in this setting). This distinction highlights the inefficiency of Flat Search in leveraging targeted traversal, which both IBis and Full Graph Search effectively utilize to process only the relevant datasets.
When varying the number of scraped datasets per model owner ($D$) (cf. Fig.10), Flat Search shows a steep increase in execution time, consistent with its dependence on $D$ in the $O(MD)$ complexity. Full Graph Search also exhibits growth with larger $D$, reflecting the added $O(D)$ term in its traversal. However, IBis maintains relatively stable execution times, as its operations focus primarily on $T$ datasets and avoid redundant processing of irrelevant datasets. This efficiency underscores the scalability of IBis compared to the other methods when dealing with large-scale datasets.
Figs.1010 show the number of model owners ($N$) and licenses per model owner ($L$). The execution times remain nearly constant for all methods. These parameters do not directly affect the graph traversal operations required for fetching models by license. The stability of IBis across these parameters further confirms its robustness and ability to handle increasing system size without compromising performance. Overall, the experimental results reaffirm the efficiency of IBis, particularly as $M$, $T$, or $D$ grows, while highlighting the limitations of Flat Search and the added overhead in Full Graph Search.